OpenStudy (anonymous):

What is the sum of the arithmetic sequence 22, 13, 4 … if there are 28 terms 1) 3,710 2) -3,094 3) -2,786 4) 4,018

OpenStudy (anonymous):

3

OpenStudy (angela210793):

y28=y1+(n-1)*d d=y2-y1=13-22=-9 y28=22+27*(-9)=22-243=-221 Sn=(y1+y28)*14=-2786 next time use the formulas :)

OpenStudy (anonymous):

i normally do the problem and then ask the question to make sure i have it right because im taking quizes and i want the best grade possible, but thanks.